# No [CII] or dust detection in two Little Red Dots at

### v1

@@
Little Red Dots (LRDs) are compact, point-like sources characterized by their red color and broad Balmer lines, which have been debated to be either dominated by active galactic nuclei (AGN) or dusty star-forming galaxies (DSFGs). Here we report two LRDs (ID9094 and ID2756) at , recently discovered in the JWST FRESCO GOODS-North field. Both satisfy the “v-shape” colors and compactness criteria for LRDs and are identified as Type-I AGN candidates based on their broad H emission lines (full width at half maximum: km s-1 for ID9094 and km s-1 for ID2756) and narrow [O iii] lines ( km s-1). To investigate their nature, we conduct deep NOEMA follow-up observations targeting the [C ii] emission line and the 1.3 mm dust continuum. We do not detect [C ii] or 1.3 mm continuum emission for either source. Notably, in the scenario that the two LRDs were DSFGs, we would expect significant detections: for [C ii] and for the 1.3 mm continuum of ID9094, and for [C ii] of ID2756. Using the upper limits of [C ii] and 1.3 mm, we perform two analyses: (1) UV-to-FIR spectral energy distribution (SED) fitting with and without AGN components, and (2) comparison of their properties with the –SFRtot empirical relation. Both analyses indicate that LRDs favor the presence of AGN to explain their observed properties. Our results highlight the importance of far-infrared observations for studying LRDs, a regime that remains largely unexplored.
@@
galaxies: high-redshift – galaxies: active – galaxies: star-formation – galaxies: photometry – submillimetre: galaxies1 Introduction
@@
Little Red Dots (LRDs; e.g., Fujimoto et al., 2022; Labbe et al., 2023; Matthee et al., 2024; Barro et al., 2024) have emerged as one of the most intriguing populations studied with the JWST (Gardner et al., 2023). The term ”LRDs” was first introduced by Matthee et al. (2024) to describe galaxies exhibiting broad H emission lines, with a red point-source morphology. In a broader context, the name LRD has also been applied to a variety of compact, red sources – predominantly at high redshifts () – with “v-shaped” (blue UV and red optical slopes in rest-frame) spectral energy distributions (SEDs), based mainly on photometric measurements only. The samples of these sources overlap, contributing to the complexity of their classification. As a result, their nature has sparked significant debate, with interpretations ranging from dusty star-forming galaxies (DSFGs; Pérez-González et al., 2024; Labbé et al., 2023; Baggen et al., 2024; Guia et al., 2024; Williams et al., 2024) to AGN-dominated systems (e.g., Kocevski et al., 2023, 2024; Ji et al., 2025), and hybrid scenarios involving both stellar and AGN contributions dominating different parts of the SED (e.g., Greene et al., 2024; Kokorev et al., 2024; Wang et al., 2024b; Akins et al., 2025).
@@
Multi-wavelength studies of LRDs have primarily focused on optical, near-infrared (NIR), mid-infrared (MIR), X-ray, and radio data, whereas far-infrared (FIR) studies remain relatively scarce in comparison. FIR observations, however, are crucial for disentangling the roles of AGN activity and dusty starburst processes, as they directly probe dust properties, star formation rates, and possible AGN contributions. Some studies have started to investigate LRDs in the FIR (e.g., Labbe et al., 2023; Casey et al., 2024; Akins et al., 2024; Williams et al., 2024), but most rely on photometric redshifts – either through stacked analyses or individual source studies – which might introduce large uncertainties in interpreting their dust properties and star formation activity. Current evidence shows that many LRDs are faint in X-rays (e.g., Ananna et al., 2024; Maiolino et al., 2024; Yue et al., 2024), exhibit weak or no radio emission (e.g., Akins et al., 2024), and lack detectable hot dust in the MIR (e.g., Pérez-González et al., 2024; Williams et al., 2024). These properties could result from heavily obscured AGNs, extreme dusty starbursts, or even hybrid scenarios. Without robust FIR constraints, however, distinguishing between these possibilities remains challenging.
@@
Another potential issue is the partial overlap in photometric selection methods for LRDs (e.g., Labbe et al., 2023; Greene et al., 2024; Kokorev et al., 2024) and optical/NIR-selected DSFGs (optically dark/faint galaxies; e.g., Wang et al., 2019; Franco et al., 2018; Alcalde Pampliega et al., 2019; Williams et al., 2019; Gómez-Guijarro et al., 2023; Xiao et al., 2023; Gómez-Guijarro et al., 2022; Barrufet et al., 2023; Williams et al., 2024; McKinney et al., 2023; Akins et al., 2023; Pérez-González et al., 2023; Barro et al., 2024), due to their similar red colors. This may pose a challenge for DSFG studies, as optical/NIR-selected DSFGs typically have more relaxed selection criteria, relying solely on color and/or magnitude cuts without considering morphology. As a result, LRDs could inadvertently be included in DSFG samples. Additionally, studies also show that some DSFGs have compact morphology (e.g., Gómez-Guijarro et al., 2023). This overlap introduces ambiguity, as both populations could appear in the same sample category despite potentially distinct physical origins. For example, DSFGs are typically powered by intense star formation obscured by large amounts of dust, while AGN-driven LRDs may exhibit similar photometric properties due to their reddened SEDs but have fundamentally different drivers of energy output.
@@
In this study, we investigate two LRDs, ID9094 and ID2756111ID9094 and ID2756 correspond to gn17537 and gn28924, respectively, in the FRESCO [O iii] catalog (Meyer et al., 2024)., located in the JWST FRESCO GOODS-North field (Oesch et al., 2023). Due to the overlap in photometric selection criteria between DSFGs and LRDs, these two sources were initially classified as optically faint, dust-obscured galaxy candidates with red colors (F182M F444W 1.5 mag) and as the most massive candidates at () based on optical-to-near-IR SED fitting (Xiao et al., 2024). However, as revealed in this study, deep JWST NIRCam grism spectra later reveal the presence of broad H emission lines (Sect. 2.4) – previously tentative in FRESCO spectra and at the time of the NOEMA proposal – along with a red, point-source-like morphology, indicating potential AGN activity.
@@
To further investigate their nature, we present follow-up FIR observations from NOEMA, targeting the [C ii] emission line and the 1.3 mm dust continuum. This study, along with Setton et al. (submitted), represents one of the first efforts to investigate LRDs with both dust and [C ii] emission. These observations provide crucial constraints on the dust temperature, allowing us to assess whether the observed properties of these LRDs are consistent with DSFGs or if AGN-driven processes dominate. By combining FIR data with existing multi-wavelength photometry, this work addresses the critical question of whether DSFG scenarios can be ruled out for LRDs in the absence of spectroscopic confirmation.
@@
This paper is organized as follows. In Section 2, we describe the data and observational details. Section 3 presents our results, including the non-detection of the dust continuum and [C ii] line in the FIR and its implications for the nature of these LRDs. In Section 4, we discuss the impact of LRDs on our understanding of DSFGs. We summarize our conclusions in Section 5.
@@
Throughout this paper, we adopt a Chabrier initial mass function (IMF; Chabrier, 2003) to estimate SFR and stellar mass. We assume a Planck cosmology (Planck Collaboration et al., 2020) with . When necessary, data from the literature have been converted with a conversion factor of (Salpeter, 1955, IMF) = 1.7 (Chabrier, 2003, IMF). All magnitudes are on the AB system (Oke & Gunn, 1983), such that log(Sν [Jy]).
@@ 2 Data and sample
The two sources ID9094 (, [J2000] = 189.019240, 62.243531) at and ID2756 (, [J2000] = 189.083488, 62.202579) at are located in the GOODS-N field, and were first discovered in the JWST FRESCO survey (Oesch et al., 2023) and reported in Xiao et al. (2024). They have a red color of F182M F444W 1.5 mag and faint F182M 26.4 mag. As such, they were selected as optically faint dusty galaxy candidates, with high inferred stellar masses. With the broad H emission line, “v-shape” colors, and compact morphology in F444W, we now confirm that they are, in fact, LRDs (see details in Sect. 2.4).
@@ 2.1 JWST observations
The two sources were initially observed by the JWST FRESCO survey (Oesch et al., 2023), which provides NIRCam/grism spectroscopy in the F444W filter and direct imaging in F182M, F210M, and F444W over 62 arcmin2 in each GOODS field, North and South. The grism spectra cover a wavelength range of 3.8 to 5.0 m at a resolution of R1,600, reaching a 5 line sensitivity of 10-18 erg s-1 cm-2. The images reach typical 5 depths of 28.3, 28.1, and 28.2 mag, respectively, in a 016 circular aperture. Data reduction was performed using the grizli pipeline222https://github.com/gbrammer/grizli, including continuum subtraction and optimal 1D spectral extraction.
Recently, we obtained deeper NIRCam/grism spectroscopy from a Cycle 3 JWST program (GO-4762; PIs: S. Fujimoto & G. Brammer), using the F410M filter instead of F444W. This choice improves spectral sensitivity by blocking the high thermal background beyond 4.4 m, at the cost of reduced wavelength coverage. Despite a slightly shorter exposure time (by a factor of 1.3) than FRESCO, this dataset thus achieves a 5 line sensitivity of 10-19 erg s-1 cm-2. The continuum subtraction was optimized to explicitly mask emission lines and a wide kernel (for the running median) to prevent over-subtraction of broad wings (see Matthee et al., 2024).
The NIRCam images and grism spectra of ID9094 and ID2756 are shown in Fig. 1, updated with the latest data.
@@
| Chandraa | 0.5–7 keV | erg s-1 cm-2 | ||
| 2–7 keV | erg s-1 cm-2 | |||
| 0.5–2 keV | erg s-1 cm-2 | |||
| HST/ACS F435W | 0.43 µm | nJy | ||
| HST/ACS F606W | 0.59 µm | nJy | ||
| HST/ACS F775W | 0.77 µm | nJy | ||
| HST/ACS F814W | 0.80 µm | nJy | ||
| HST/ACS F850LP | 0.90 µm | nJy | ||
| HST/WFC3 F105W | 1.06 µm | nJy | 33.0 8.4 | … |
| HST/WFC3 F125W | 1.25 µm | nJy | 87.5 10.8 | 75.7 9.8 |
| HST/WFC3 F140W | 1.40 µm | nJy | 84.7 6.6 | … |
| HST/WFC3 F160W | 1.54 µm | nJy | 91.3 7.8 | 109.7 10.4 |
| JWST/NIRCam F090W | 0.90 µm | nJy | ||
| JWST/NIRCam F115W | 1.15 µm | nJy | 79.2 4.4 | 64.8 6.2 |
| JWST/NIRCam F150W | 1.50 µm | nJy | … | 93.7 12.9 |
| JWST/NIRCam F182M | 1.85 µm | nJy | 80.0 4.0 | 99.3 6.7 |
| JWST/NIRCam F200W | 1.99 µm | nJy | … | 109.6 11.5 |
| JWST/NIRCam F210M | 2.10 µm | nJy | 117.8 5.9 | 130.2 8.8 |
| JWST/NIRCam F356W | 3.57 µm | nJy | 225.8 11.3 | 443.8 22.2 |
| JWST/NIRCam F444W | 4.40 µm | nJy | 372.4 18.6 | 887.7 44.4 |
| Spitzer/IRAC | 3.6 µm | nJy | 191.0 15.4 | 400.4 13.0 |
| 4.5 µm | nJy | 340.2 17.6 | 592.7 16.5 | |
| 5.8 µm | nJy | 933.7 243.5 | ||
| 8.0 µm | nJy | 351.2 284.8 | 2683.0 289.4 | |
| Spitzer/MIPSb | 24 µm | µJy | ||
| 70 µm | mJy | |||
| Herschel/PACSb | 100 µm | mJy | ||
| 160 µm | mJy | |||
| Herschel/SPIREb | 250 µm | mJy | ||
| 350 µm | mJy | |||
| 500 µm | mJy | |||
| SCUBA-2c | 850 µm | mJy | ||
| NOEMA | 1.3 mm | µJy | ||
| [C ii]158μm | Jy km s-1 | |||
| NIKA-2d | 1.2 mm | µJy | ||
| 2.0 mm | µJy | |||
| VLA 1.5 GHze | 20 cm | uJy |
Note: We use 1 for uncertainties and 3 upper limits for non-detections. (a) We calculate 3 upper limits based on the mean sensitivity limits achieved in the central 1 arcmin2 area at the average aim point from 2 Ms Chandra Deep Field-North survey (Xue et al., 2016). (b) Sources are not detected, we adopt 3 upper limits from the GOODS-Herschel survey (Elbaz et al., 2011). (c) We adopt the confusion limit for a 3 detection in the central region of GOODS-N from Cowie et al. (2017). (d) We adopt 3 upper limits according to the N2CLS depth from Bing et al. (2023). (e) Calculated based on the rms noise for a 16 resolution in the field center from Owen (2018).
| ID | FluxHβ,broad | FluxHβ,narrow | Flux | |||
| erg s-1cm-2 | erg s-1cm-2 | erg s-1cm-2 | km s-1 | km s-1 | ||
| ID9094 | ||||||

### v2

@@
Little Red Dots (LRDs) are compact, point-like sources characterized by their red color and broad Balmer lines, which have been debated to be either dominated by active galactic nuclei (AGN) or dusty star-forming galaxies (DSFGs). Here we report two LRDs (ID9094 and ID2756) at , recently discovered in the JWST FRESCO GOODS-North field. Both satisfy the āv-shapeā colors and compactness criteria for LRDs and are identified as Type-I AGN candidates based on their broad H emission lines (full width at half maximum: km s-1 for ID9094 and km s-1 for ID2756) and narrow [Oāiii]ā lines ( km s-1). To investigate their nature, we conduct deep NOEMA follow-up observations targeting the [Cāii] emission line and the 1.3 mm dust continuum. We do not detect [Cāii]ā or 1.3 mm continuum emission for either source. Notably, in the scenario that the two LRDs were DSFGs, we would expect significant detections: for [Cāii]Ā and for the 1.3 mm continuum of ID9094, and for [Cāii]Ā of ID2756. Using the upper limits of [Cāii]ā and 1.3 mm, we perform two analyses: (1) UV-to-FIR spectral energy distribution (SED) fitting with and without AGN components, and (2) comparison of their properties with the āSFRtot empirical relation. Both analyses are consistent with a scenario where AGN activity may contribute to the observed properties, though a dusty star-forming origin cannot be fully ruled out. Our results highlight the importance of far-infrared observations for studying LRDs, a regime that remains largely unexplored.
@@
galaxies: high-redshift ā galaxies: active ā galaxies: star-formation ā galaxies: photometry ā submillimetre: galaxies1 Introduction
@@
Little Red Dots (LRDs; e.g., Fujimoto etĀ al., 2022; Labbe etĀ al., 2025; Matthee etĀ al., 2024; Barro etĀ al., 2024) have emerged as one of the most intriguing populations studied with the JWST (Gardner etĀ al., 2023). The term āLRDsā was first introduced by Matthee etĀ al. (2024) to describe galaxies exhibiting broad H emission lines, with a red point-source morphology. In a broader context, the name LRD has also been applied to a variety of compact, red sources ā predominantly at high redshifts () ā with āv-shapedā (blue UV and red optical slopes in rest-frame) spectral energy distributions (SEDs), based mainly on photometric measurements only. The samples of these sources overlap, contributing to the complexity of their classification. Most recent studies agree that LRDs involve a combination of stellar and AGN contributions, although there is ongoing debate over which component dominates (e.g., Wang etĀ al., 2024; Akins etĀ al., 2025a). Some works emphasize a dusty star-forming galaxy-dominated interpretation (e.g., PĆ©rez-GonzĆ”lez etĀ al., 2024; LabbĆ© etĀ al., 2023; Baggen etĀ al., 2024; Williams etĀ al., 2024), while others find stronger evidence for AGN dominance (e.g., Kocevski etĀ al., 2023, 2025; Greene etĀ al., 2024; Kokorev etĀ al., 2024; Ji etĀ al., 2025; Naidu etĀ al., 2025; de Graaff etĀ al., 2025).
@@
Multi-wavelength studies of LRDs have primarily focused on rest-frame UV, optical, near-infrared (NIR), mid-infrared (MIR), X-ray, and radio data, whereas far-infrared (FIR) studies remain relatively scarce in comparison. FIR observations, however, are crucial for disentangling the roles of AGN activity and dusty starburst processes, as they directly probe dust properties, star formation rates, and possible AGN contributions. Some studies have started to investigate LRDs in the FIR (e.g., Labbe etĀ al., 2025; Casey etĀ al., 2024, 2025; Akins etĀ al., 2024, 2025b; Williams etĀ al., 2024; Setton etĀ al., 2025), but most rely on photometric redshifts ā either through stacked analyses or individual source studies ā which might introduce large uncertainties in interpreting their dust properties and star formation activity. Current evidence shows that many LRDs are faint in X-rays (e.g., Ananna etĀ al., 2024; Maiolino etĀ al., 2025; Yue etĀ al., 2024; Sacchi & Bogdan, 2025), exhibit weak or no radio emission (e.g., Akins etĀ al., 2024), and lack detectable hot dust in the MIR (e.g., PĆ©rez-GonzĆ”lez etĀ al., 2024; Williams etĀ al., 2024). These properties could result from heavily obscured AGNs, extreme dusty starbursts, or even hybrid scenarios. Without robust FIR constraints, however, distinguishing between these possibilities remains challenging.
@@
Another potential issue is the partial overlap in photometric selection methods for LRDs (e.g., Labbe et al., 2025; Greene et al., 2024; Kokorev et al., 2024) and optical/NIR-selected DSFGs (optically dark/faint galaxies; e.g., Wang et al., 2019; Franco et al., 2018; Alcalde Pampliega et al., 2019; Williams et al., 2019; Gómez-Guijarro et al., 2023; Xiao et al., 2023; Gómez-Guijarro et al., 2022; Barrufet et al., 2023; Williams et al., 2024; McKinney et al., 2023; Akins et al., 2023; Pérez-GonzÔlez et al., 2023; Barro et al., 2024), due to their similar red colors. This may pose a challenge for DSFG studies, as optical/NIR-selected DSFGs typically have more relaxed selection criteria, relying solely on color and/or magnitude cuts without considering morphology. As a result, LRDs could inadvertently be included in DSFG samples. Additionally, studies also show that some DSFGs have compact morphology (e.g., Gómez-Guijarro et al., 2023). This overlap introduces ambiguity, as both populations could appear in the same sample category despite potentially distinct physical origins. For example, DSFGs are typically powered by intense star formation obscured by large amounts of dust, while AGN-driven LRDs may exhibit similar photometric properties due to their reddened SEDs but have fundamentally different drivers of energy output.
@@
In this study, we investigate two LRDs, ID9094 and ID2756111ID9094 and ID2756 correspond to gn17537 and gn28924, respectively, in the FRESCO [Oāiii]ā catalog (Meyer etĀ al., 2024)., located in the JWST FRESCO GOODS-North field (Oesch etĀ al., 2023). Due to the overlap in photometric selection criteria between DSFGs and LRDs, these two sources were initially classified as optically faint, dust-obscured galaxy candidates with red colors (F182M F444W 1.5 magnitude; hereafter mag) and as the most massive candidates at () based on optical-to-near-IR SED fitting (Xiao etĀ al., 2024). However, as revealed in this study, deep JWST NIRCam grism spectra later reveal the presence of broad H emission lines (broad component of with full width at half maximum 1000 km s-1; e.g., Matthee etĀ al. (2024); Sect.Ā 2.4)222The broad lines were not clearly identified in the initial FRESCO spectra or at the time of the NOEMA proposal, due to the limited sensitivity of the FRESCO data., along with a red, point-source-like morphology, indicating potential AGN activity.
@@
To further investigate their nature, we present follow-up FIR observations from NOEMA, targeting the [Cāii] emission line and the 1.3 mm dust continuum. This study, together with the recent work by Setton etĀ al. (2025), represents one of the first efforts to investigate LRDs with both dust and [Cāii]ā emission. Setton etĀ al. (2025) focused on two of the most luminous LRDs at and reported a deficit of hot and cold dust emission. In comparison, our study targets two LRDs at , a redshift regime more representative of the typical LRD population identified in JWST surveys. Together, the two studies provide a more complete view of the dust and gas properties of LRDs. Our NOEMA observations provide crucial constraints on the dust temperature, allowing us to assess whether the observed properties of these LRDs are consistent with DSFGs or if AGN-driven processes dominate. By combining FIR data with existing multi-wavelength photometry, this work addresses the critical question of whether DSFG scenarios can be ruled out for LRDs in the absence of spectroscopic confirmation.
@@
This paper is organized as follows. In SectionĀ 2, we describe the data and observational details. SectionĀ 3 presents our results, including the non-detection of the dust continuum and [Cāii]āline in the FIR and its implications for the nature of these LRDs. In SectionĀ 4, we discuss the impact of LRDs on our understanding of DSFGs. We summarize our conclusions in SectionĀ 5.
@@
Throughout this paper, we adopt a Chabrier initial mass function (IMF; Chabrier, 2003) to estimate SFR and stellar mass. We assume a Planck cosmology (Planck Collaboration et al., 2020) with . When necessary, data from the literature have been converted with a conversion factor of (Salpeter, 1955, IMF) = 1.7 (Chabrier, 2003, IMF). All magnitudes are in the AB system (Oke & Gunn, 1983), such that log(Sν [Jy]).
@@ 2 Data and sample
The two sources ID9094 (, [J2000] = 189.019240, 62.243531) at and ID2756 (, [J2000] = 189.083488, 62.202579) at are located in the GOODS-N field, and were first discovered in the JWST FRESCO survey (Oesch etĀ al., 2023) and reported in Xiao etĀ al. (2024). They have red colors (F182M F444W 1.5 mag) and are faint in the F182M band (F182M 26.4 mag). As such, they were selected as optically faint dusty galaxy candidates, with high inferred stellar masses. In this study, based on their broad H emission line, āv-shapeā colors, and compact morphology in F444W, we now identify both sources as LRDs, following commonly used criteria (e.g., Matthee etĀ al. 2024; Labbe etĀ al. 2025; Greene etĀ al. 2024; Kokorev etĀ al. 2024; Kocevski etĀ al. 2025; see details in Sect.Ā 2.4).
@@ 2.1 JWST observations
The two sources were initially observed by the JWST FRESCO survey (Oesch etĀ al., 2023), which provides NIRCam/grism spectroscopy in the F444W filter and direct imaging in F182M, F210M, and F444W over 62 arcmin2 in each GOODS field, North and South. The grism spectra cover a wavelength range of 3.8 to 5.0 m at a resolution of R1,600, reaching a 5 line sensitivity of 10-18 ergās-1ācm-2. The images reach typical 5 depths of 28.3, 28.1, and 28.2 mag, respectively, in a 016 circular aperture. Data reduction was performed using the grizli pipeline333https://github.com/gbrammer/grizli, including continuum subtraction and optimal 1D spectral extraction.
Recently, we obtained deeper NIRCam/grism spectroscopy from a Cycle 3 JWST program (GO-4762; PIs: S. Fujimoto & G.Ā Brammer), using the F410M filter instead of F444W. This choice improves spectral sensitivity by blocking the high thermal background beyond 4.4 m, at the cost of reduced wavelength coverage. Despite a slightly shorter exposure time (by a factor of 1.3) than FRESCO, this dataset thus achieves a 5 line sensitivity of 10-19 ergās-1ācm-2. The continuum subtraction was optimized to explicitly mask emission lines and a wide kernel with a large central hole (151, 31 pixels, respectively; for the running median) to prevent over-subtraction of broad wings (see Matthee etĀ al., 2024).
The NIRCam images and grism spectra of ID9094 and ID2756 are shown in Fig.Ā 1, updated with the latest data.
@@
| Chandraa | 0.5ā7 keV | ergās-1ācm-2 | ||
| 2ā7 keV | ergās-1ācm-2 | |||
| 0.5ā2 keV | ergās-1ācm-2 | |||
| HST/ACS F435W | 0.43āµm | nJy | ||
| HST/ACS F606W | 0.59āµm | nJy | ||
| HST/ACS F775W | 0.77āµm | nJy | ||
| HST/ACS F814W | 0.80āµm | nJy | ||
| HST/ACS F850LP | 0.90āµm | nJy | ||
| HST/WFC3 F105W | 1.06āµm | nJy | 33.0 8.4 | ⦠|
| HST/WFC3 F125W | 1.25āµm | nJy | 87.5 10.8 | 75.7 9.8 |
| HST/WFC3 F140W | 1.40āµm | nJy | 84.7 6.6 | ⦠|
| HST/WFC3 F160W | 1.54āµm | nJy | 91.3 7.8 | 109.7 10.4 |
| JWST/NIRCam F090W | 0.90āµm | nJy | ||
| JWST/NIRCam F115W | 1.15āµm | nJy | 79.2 4.4 | 64.8 6.2 |
| JWST/NIRCam F150W | 1.50āµm | nJy | ⦠| 93.7 12.9 |
| JWST/NIRCam F182M | 1.85āµm | nJy | 80.0 4.0 | 99.3 6.7 |
| JWST/NIRCam F200W | 1.99āµm | nJy | ⦠| 109.6 11.5 |
| JWST/NIRCam F210M | 2.10āµm | nJy | 117.8 5.9 | 130.2 8.8 |
| JWST/NIRCam F356W | 3.57āµm | nJy | 225.8 11.3 | 443.8 22.2 |
| JWST/NIRCam F444W | 4.40āµm | nJy | 372.4 18.6 | 887.7 44.4 |
| Spitzer/IRAC | 3.6āµm | nJy | 191.0 15.4 | 400.4 13.0 |
| 4.5āµm | nJy | 340.2 17.6 | 592.7 16.5 | |
| 5.8āµm | nJy | 933.7 243.5 | ||
| 8.0āµm | nJy | 351.2 284.8 | 2683.0 289.4 | |
| Spitzer/MIPSb | 24āµm | µJy | ||
| 70āµm | mJy | |||
| Herschel/PACSb | 100āµm | mJy | ||
| 160āµm | mJy | |||
| Herschel/SPIREb | 250āµm | mJy | ||
| 350āµm | mJy | |||
| 500āµm | mJy | |||
| SCUBA-2c | 850āµm | mJy | ||
| NOEMA | 1.3āmm | µJy | ||
| [Cāii]158μm | Jy km s-1 | |||
| NIKA-2d | 1.2āmm | µJy | ||
| 2.0āmm | µJy | |||
| VLA 1.5āGHze | 20ācm | uJy |
ā¢
Note: We use 1 for uncertainties and 3 upper limits for non-detections. (a) We calculate 3 upper limits based on the mean sensitivity limits achieved in the central 1 arcmin2 area at the average aim point from 2 Ms Chandra Deep Field-North survey (Xue etĀ al., 2016). (b) Sources are not detected, we adopt 3 upper limits from the GOODS-Herschel survey (Elbaz etĀ al., 2011). (c) We adopt the confusion limit for a 3 detection in the central region of GOODS-N from Cowie etĀ al. (2017). (d) We adopt 3 upper limits according to the N2CLS depth from Bing etĀ al. (2023). (e) Calculated based on the rms noise for a 16 resolution in the field center from Owen (2018).
| ID | FluxHβ,broad | FluxHβ,narrow | Flux | |||||
| erg s-1cm-2 | erg s-1cm-2 | erg s-1cm-2 | km s-1 | km s-1 | ||||
| ID9094 | ||||||||
@@
ā¢
Note. and are continuum slopes derived from multi-band JWST photometry blueward and redward of the Balmer break, respectively, and trace the two sides of the āv-shapeā SEDs (see Sect.Ā 2.4).